# Improved Noise Schedule for Diffusion Training

### v1

@@ Abstract
Diffusion models have emerged as the de facto choice for generating visual signals. However, training a single model to predict noise across various levels poses significant challenges, necessitating numerous iterations and incurring significant computational costs. Various approaches, such as loss weighting strategy design and architectural refinements, have been introduced to expedite convergence. In this study, we propose a novel approach to design the noise schedule for enhancing the training of diffusion models. Our key insight is that the importance sampling of the logarithm of the Signal-to-Noise ratio ($\log\text{SNR}$), theoretically equivalent to a modified noise schedule, is particularly beneficial for training efficiency when increasing the sample frequency around $\log\text{SNR}=0$. We empirically demonstrate the superiority of our noise schedule over the standard cosine schedule. Furthermore, we highlight the advantages of our noise schedule design on the ImageNet benchmark, showing that the designed schedule consistently benefits different prediction targets.
@@ 1 Introduction
Diffusion models have emerged as a pivotal technique for generating visual signals across diverse domains, such as image synthesis Ramesh et al. (2022); Saharia et al. (2022); Rombach et al. (2022) and video generation Brooks et al. (2024). They are particularly adept at approximating complex distributions, where Generative Adversarial Networks (GANs) may encounter difficulties. Despite the substantial computational resources and numerous training iterations required for convergence, improving the training efficiency of diffusion models is essential for their application in large-scale scenarios, such as high-resolution image synthesis and video generation.
Architectural enhancements offer a promising path to improve both the training speed and performance of diffusion models. For instance, the use of Adaptive Layer Normalization Gu et al. (2022), when combined with zero initialization in the Transformer architecture as demonstrated by Peebles & Xie (2023), represents such an improvement. Similarly, the adoption of U-shaped skip connections within Transformers, as outlined in previous works Hoogeboom et al. (2023); Bao et al. (2022); Crowson et al. (2024), also boosts efficiency. In a parallel development, Karras et al. (2024) have contributed to this endeavor by reengineering the layers of ADM UNet Dhariwal & Nichol (2021) to preserve the magnitudes of activations, weights, and updates, ensuring a more efficient learning process.
Concurrently, various loss weighting designs have been implemented to accelerate the convergence of training. Previous works, such as eDiff-I Balaji et al. (2022) and Min-SNR Hang et al. (2023), found that the training of diffusion models may encounter conflicts among various noise intensities. Choi et al. (2022) prioritize specific noise levels during training to enhance learning of visual concepts. Min-SNR Hang et al. (2023) reduces weights of noisy tasks, pursuing the Pareto Optimality in different denoising tasks, validated its effectiveness on multiple datasets and architectures. A softer version of this approach, aiming to further enhance high-resolution image synthesis within hourglass diffusion models, was introduced by Crowson et al. (2024). SD3 Esser et al. (2024) empirically found that it’s crucial to increase the weight of the intermediate noise intensities, which has demonstrated the effectiveness during training the diffusion models.
In this study, we present a novel method to enhance the training of diffusion models by strategically redefining the noise schedule, which is equivalent to importance sampling of the noise across different intensities. However, empirical evidence suggests that allocating more computation costs (FLOPs) to mid-range noise levels (around $\log\text{SNR}=0$) yields superior performance compared to increasing loss weights during the same period, particularly under constrained computational budgets. We experimentally analyze the performance of several different noise schedules, including Laplace, Cauchy, and the Cosine Shifted/Scaled, which are visualized in Figure 1. Notably, the Laplace schedule exhibits favorable performance. We recommend to choose this noise schedule in the future.
We demonstrate the effectiveness of this approach using the ImageNet benchmark, with a consistent training budget of 500K iterations. Evaluated using the FID metric, our results reveal that noise schedules with a concentrated probability density around $\log\text{SNR}=0$ consistently surpass others, as evidenced at both $256\times 256$ and $512\times 512$ resolutions with different prediction target. This research contributes to the advancement of efficient training techniques for diffusion models.
@@ 2.1 Preliminaries
Diffusion models Ho et al. (2020); Yang et al. (2021) learn to generate data by iteratively reversing the diffusion process. We denote the distribution of data points as $\mathbf{x}\sim p_{\text{data}}(\mathbf{x})$. The diffusion process progressively adds noise to the data, which is defined as:
@@
where $\alpha_{t}$ and $\sigma_{t}$ are the coefficients of the adding noise process, essentially representing the noise schedule. For the commonly used prediction target velocity: $\mathbf{v}_{t}=\alpha_{t}\bm{\epsilon}-\sigma_{t}\mathbf{x}$ Salimans & Ho (2022), the diffusion model $\theta$ is trained through the Mean Squared Error (MSE) loss:
@@
where $w(t)$ is the loss weight, $\mathbf{c}$ denotes the condition information. Common practices sample $t$ from the uniform distribution $\mathcal{U}[0,1]$. Kingma et al. (2021) introduced the Signal-to-Noise ratio as $\text{SNR}(t)=\frac{\alpha_{t}^{2}}{\sigma_{t}^{2}}$ to measure the noise level of different states. To simplify, we denote $\lambda=\log\text{SNR}$ to indicate the noise intensities. In the Variance Preserving (VP) setting, the coefficients in Equation 1 can be calculated by $\alpha^{2}_{t}=\frac{\exp(\lambda)}{\exp(\lambda)+1}$, $\sigma^{2}_{t}=\frac{1}{\exp(\lambda)+1}$.
2.2 Improved Noise Schedule Design
Given that the timestep $t$ is a random variable sampled from uniform distribution, the noise schedule implicitly defines the distribution of importance sampling on various noise levels. The sampling probability of noise intensity $\lambda$ is:
$p(\lambda)=p(t)\left|\frac{\mathrm{d}t}{\mathrm{d}\lambda}\right|.$ (3)
Considering that $t$ satisfies uniform distribution, and $\lambda$ is monotonically decreasing with $t$, we have:
$p(\lambda)=-\frac{\mathrm{d}t}{\mathrm{d}\lambda}.$ (4)
We take cosine noise schedule Nichol & Dhariwal (2021) as an example, where $\alpha_{t}=\cos\left(\frac{\pi t}{2}\right)$, $\sigma_{t}=\sin\left(\frac{\pi t}{2}\right)$. Then we can deduce that $\lambda=-2\log\tan(\pi t/2)$ and $t=2/\pi\arctan e^{-\lambda/2}$. Thus the distribution of $\lambda$ is: $p(\lambda)=-\mathrm{d}t/\mathrm{d}\lambda=\text{sech}(\lambda/2)/2\pi$. This derivation illustrates the process of obtaining $p(\lambda)$ from a noise schedule $\lambda(t)$. On the other hand, we can derive the noise schedule from the sampling probability of different noise intensities $p(\lambda)$. By integrating Equation 4, we have:
@@ 2.2 Improved Noise Schedule Design
where $\mathcal{P}(\lambda)$ represents the cumulative distribution function of $\lambda$. Thus we can obtain the noise schedule $\lambda$ by applying the inverse function $\mathcal{P}^{-1}$. In conclusion, during the training process, the importance sampling of varying noise intensities essentially equates to the modification of the noise schedules.
@@
Cosine $\text{sech}\left(\lambda/2\right)/2\pi$ $2\log\cot\left(\frac{\pi t}{2}\right)$
@@
Table 1: Overview of various Noise Schedules. The table categorizes them into five distinct types: Cosine, Laplace, Cauchy, and two variations of Cosine schedules. The second column $p(\lambda)$ denotes the sampling probability at different noise intensities $\lambda$. The last column $\lambda(t)$ indicates how to sample noise intensities for training. We derived their relationship in Equation 4 and Equation 6.
@@ 2.3 Unified Formulation for Diffusion Training
where $\mathcal{D}$ signifies the training dataset, noise $\bm{\epsilon}$ is drawn from a standard Gaussian distribution, and $p(\lambda)$ is the distribution of noise intensities. Different predicting targets, such as $\mathbf{x}_{0}$ and $\mathbf{v}$, can also be re-parameterized to $\bm{\epsilon}$-prediction. $w(\lambda)$ denotes the loss weighting strategy. Although adjusting $w(\lambda)$ is theoretically equivalent to altering $p(\lambda)$. In practical training, directly modifying $p(\lambda)$ to concentrate computational resources on training specific noise levels is more effective than enlarging the loss weight on specific noise levels. Therefore, we focus on how to design $p(\lambda)$.
@@ 2.4 Practical Settings
By re-allocating the computation resources at different noise intensities, we can train the complete denoising process. During sampling process, we standardize the sampled SNR to align with the cosine schedule, thereby focusing our exploration solely on the impact of different strategies during training. It is important to note that, from the perspective of the noise schedule, how to allocate the computation resource during inference is also worth reconsideration. We will not explore it in this paper and leave this as future work.
@@ 3 Experiments
Dataset. We conduct experiments on ImageNet Deng et al. (2009) with $256\times 256$ and $512\times 512$ resolution. For each image, we follow the preprocessing in Rombach et al. (2022) to center crop and encode images to latents. The shape of compressed latent feature is $32\times 32\times 4$ for $256^{2}$ images and $64\times 64\times 4$ for $512^{2}$ images.
Network Architecture. We adopt DiT-B from Peebles & Xie (2023) as our backbone. We replace the last AdaLN Linear layer with vanilla linear. Others are kept the same as the original implementation.
Training Settings. We adopt the Adam optimizer with learning rate $1\times 10^{-4}$. We set the batch size to 256 as Peebles & Xie (2023); Hang et al. (2023). Each model is trained for 500K iterations if not specified. Our implementation is mainly built on OpenDiT Zhao et al. (2024) and experiments are mainly conducted on 8$\times$16G V100 GPUs.
Baselines and Metrics. We compare our proposed noise schedule with several baseline settings in Table 2. For each setting, we sample images using DDIM Song et al. (2021) with 50 steps. Despite the noise strategy for different settings may be different, we ensure they are the same at each sampling step. This approach is adopted to exclusively investigate the impact of the noise strategy during the training phase. Moreover, we report results with different classifier-free guidance scales, and the FID is calculated using 10K generated images.
@@
Table 2: Comparison of different methods and related loss weighting strategies. The $w(\lambda)$ is introduced in Equation 7.
3.2 Comparison with baselines and loss weight designs

### v2

@@ Abstract
Diffusion models have emerged as the de facto choice for generating high-quality visual signals across various domains. However, training a single model to predict noise across various levels poses significant challenges, necessitating numerous iterations and incurring significant computational costs. Various approaches, such as loss weighting strategy design and architectural refinements, have been introduced to expedite convergence and improve model performance. In this study, we propose a novel approach to design the noise schedule for enhancing the training of diffusion models. Our key insight is that the importance sampling of the logarithm of the Signal-to-Noise ratio ($\log\text{SNR}$), theoretically equivalent to a modified noise schedule, is particularly beneficial for training efficiency when increasing the sample frequency around $\log\text{SNR}=0$. This strategic sampling allows the model to focus on the critical transition point between signal dominance and noise dominance, potentially leading to more robust and accurate predictions. We empirically demonstrate the superiority of our noise schedule over the standard cosine schedule. Furthermore, we highlight the advantages of our noise schedule design on the ImageNet benchmark, showing that the designed schedule consistently benefits different prediction targets. Our findings contribute to the ongoing efforts to optimize diffusion models, potentially paving the way for more efficient and effective training paradigms in the field of generative AI.
@@ 1 Introduction
Diffusion models have emerged as a pivotal technique for generating high-quality visual signals across diverse domains, including image synthesis Ramesh et al. (2022); Saharia et al. (2022); Rombach et al. (2022) , video generation Ho et al. (2022); Singer et al. (2023); Brooks et al. (2024), and even 3D object generation Wang et al. (2022); Nichol et al. (2022). One of the key strengths of diffusion models lies in their ability to approximate complex distributions, where Generative Adversarial Networks (GANs) may encounter difficulties. Despite the substantial computational resources and numerous training iterations required for convergence, improving the training efficiency of diffusion models is essential for their application in large-scale scenarios, such as high-resolution image synthesis and long video generation.
Recent efforts to enhance diffusion model training efficiency have primarily focused on two directions. The first approach centers on architectural improvements. For instance, the use of Adaptive Layer Normalization Gu et al. (2022), when combined with zero initialization in the Transformer architecture Peebles & Xie (2023), has shown promising results. MM-DiT Esser et al. (2024) extends this approach to multi-modality by employing separate weights for vision and text processing. Similarly, U-shaped skip connections within Transformers Hoogeboom et al. (2023); Bao et al. (2022); Crowson et al. (2024) and reengineered layer designs Karras et al. (2024) have contributed to more efficient learning processes.
The second direction explores various loss weighting strategies to accelerate training convergence. Works such as eDiff-I Balaji et al. (2022) and Ernie-ViLG 2.0 Feng et al. (2022) address training difficulties across noise intensities using a Mixture of Experts approach. Other studies have investigated prioritizing specific noise levels Choi et al. (2022) and reducing weights of noisy tasks Hang et al. (2023) to enhance learning effectiveness. Recent developments include a softer weighting approach for high-resolution image synthesis Crowson et al. (2024) and empirical findings on the importance of intermediate noise intensities Esser et al. (2024).
Despite these advances, the fundamental role of noise scheduling in diffusion model training remains underexplored. In this study, we present a novel approach focusing on the fundamental role of noise scheduling, which is a function that determines how much noise is added to the input data at each timestep $t$ during the training process, controlling the distribution of noise levels that the neural network learns to remove. Our framework provides a unified perspective for analyzing noise schedules and importance sampling, leading to a straightforward method for designing noise schedules through the identification of curves in the $p(\lambda)$ distribution, as visualized in Figure 1. Through empirical analysis, we discover that allocating more computation costs (FLOPs) to mid-range noise levels (around $\log\text{SNR}=0$) yields superior performance compared to increasing loss weights during the same period, particularly under constrained computational budgets.
We evaluate several different noise schedules, including Laplace, Cauchy, and the Cosine Shifted/Scaled variants, through comprehensive experiments using the ImageNet benchmark with a consistent training budget of 500K iterations (about 100 epochs). Our results, measured using the Fréchet Inception Distance (FID) metric at both $256\times 256$ and $512\times 512$ resolutions, demonstrate that noise schedules with concentrated probability density around $\log\text{SNR}=0$ consistently outperform alternatives, with the Laplace schedule showing particularly favorable performance.
The key contributions of our work can be summarized as follows:
  • A unified framework for analyzing and designing noise schedules in diffusion models, offering a more systematic approach to noise schedule optimization.
  • Empirical evidence demonstrating the superiority of mid-range noise level focus over loss weight adjustments for improving training efficiency.
  • Comprehensive evaluation and comparison of various noise schedules, providing practical guidelines for future research and applications in diffusion model training.
@@ 2.1 Preliminaries
Diffusion models Ho et al. (2020); Yang et al. (2021) learn to generate data by iteratively reversing the diffusion process. We denote the distribution of data points as $\mathbf{x}\sim p_{\text{data}}(\mathbf{x})$. The diffusion process systematically introduces noise to the data in a progressive manner. In a continuous setting, the noisy data at timestep $t$ is defined as follows:
@@
where $\alpha_{t}$ and $\sigma_{t}$ are the coefficients of the adding noise process, essentially representing the noise schedule. For the commonly used prediction target velocity: $\mathbf{v}_{t}=\alpha_{t}\bm{\epsilon}-\sigma_{t}\mathbf{x}$ Salimans & Ho (2022), the diffusion model $\mathbf{v}_{\theta}$ is trained through the Mean Squared Error (MSE) loss:
@@
where $w(t)$ is the loss weight, $\mathbf{c}$ denotes the condition information. In the context of class-conditional generation tasks, $\mathbf{c}$ represents the class label. Common practices sample $t$ from the uniform distribution $\mathcal{U}[0,1]$. Kingma et al. (2021) introduced the Signal-to-Noise ratio as $\text{SNR}(t)=\frac{\alpha_{t}^{2}}{\sigma_{t}^{2}}$ to measure the noise level of different states. Notably, $\text{SNR}(t)$ monotonically decreases with increasing $t$. Some works represent the loss weight from the perspective of SNR Salimans & Ho (2022); Hang et al. (2023); Crowson et al. (2024). To simplify, we denote $\lambda=\log\text{SNR}$ to indicate the noise intensities. In the Variance Preserving (VP) setting, the coefficients in Equation 1 can be calculated by $\alpha^{2}_{t}=\frac{\exp(\lambda)}{\exp(\lambda)+1}$, $\sigma^{2}_{t}=\frac{1}{\exp(\lambda)+1}$.
While these foundational concepts have enabled significant progress in diffusion models, the choice of noise schedule remains somewhat ad hoc. This motivates us to develop a more systematic framework for analyzing and designing noise schedules by examining them from a probability perspective.
2.2 Noise Schedule Design from A Probability Perspective
The training process of diffusion models involves sampling timesteps $t$ from a uniform distribution. However, this uniform sampling in time actually implies a non-uniform sampling of noise intensities. We can formalize this relationship through the lens of importance sampling Bishop & Nasrabadi (2006). Specifically, when $t$ follows a uniform distribution, the sampling probability of noise intensity $\lambda$ is given by:
$p(\lambda)=p(t)\left|\frac{\mathrm{d}t}{\mathrm{d}\lambda}\right|=-\frac{\mathrm{d}t}{\mathrm{d}\lambda},$ (3)
where the negative sign appears because $\lambda$ monotonically decreases with $t$. We take cosine noise schedule Nichol & Dhariwal (2021) as an example, where $\alpha_{t}=\cos\left(\frac{\pi t}{2}\right)$, $\sigma_{t}=\sin\left(\frac{\pi t}{2}\right)$. Then we can deduce that $\lambda=-2\log\tan(\pi t/2)$ and $t=2/\pi\arctan e^{-\lambda/2}$. Thus the distribution of $\lambda$ is: $p(\lambda)=-\mathrm{d}t/\mathrm{d}\lambda=\text{sech}(\lambda/2)/2\pi$. This derivation illustrates the process of obtaining $p(\lambda)$ from a noise schedule $\lambda(t)$. On the other hand, we can derive the noise schedule from the sampling probability of different noise intensities $p(\lambda)$. By integrating Equation 3, we have:
@@ 2.2 Noise Schedule Design from A Probability Perspective
where $\mathcal{P}(\lambda)$ represents the cumulative distribution function of $\lambda$. Thus we can obtain the noise schedule $\lambda$ by applying the inverse function $\mathcal{P}^{-1}$. In conclusion, during the training process, the importance sampling of varying noise intensities essentially equates to the modification of the noise schedules. To illustrate this concept, let’s consider the Laplace distribution as an example , we can derive the cumulative distribution function $\mathcal{P}(\lambda)=1-\int\frac{1}{2b}\exp\left(-|\lambda-\mu|/{b}\right)\mathrm{d}\lambda=\frac{1}{2}\left(1+\text{sgn}(\lambda-\mu)(1-\exp(-|\lambda-\mu|/b))\right)$. Subsequently, we can obtain the inverse function to express the noise schedule in terms of $\lambda$: $\lambda=\mu-b\text{sgn}(0.5-t)\ln(1-2|t-0.5|)$. Here, $\text{sgn}(\cdot)$ denotes the signum function, which equals 1 for positive inputs, $-1$ for negative inputs. The pseudo-code for implementing the Laplace schedule in the training of diffusion models is presented in A.1.
@@
Cosine $\text{sech}\left(\lambda/2\right)/2\pi$ $2\log\left(\cot\left(\frac{\pi t}{2}\right)\right)$
@@
Table 1: Overview of various Noise Schedules. The table categorizes them into five distinct types: Cosine, Laplace, Cauchy, and two variations of Cosine schedules. The second column $p(\lambda)$ denotes the sampling probability at different noise intensities $\lambda$. The last column $\lambda(t)$ indicates how to sample noise intensities for training. We derived their relationship in Equation 3 and 5.
This framework reveals that noise schedule design can be reframed as a probability distribution design problem. Rather than directly specifying how noise varies with time, we can instead focus on how to optimally distribute our sampling across different noise intensities. Our approach is also applicable to the recently popular flow matching with logit normal sampling scheme Esser et al. (2024). Within our framework, we analyzed the distribution of its logSNR in A.4 and demonstrated its superiority over vanilla flow matching and cosine scheduling from the perspective of $p(\lambda)$.
@@ 2.3 Unified Formulation for Diffusion Training
where $\mathcal{D}$ signifies the training dataset, noise $\bm{\epsilon}$ is drawn from a standard Gaussian distribution, and $p(\lambda)$ is the distribution of noise intensities. This formulation provides a flexible framework that can accommodate various diffusion training strategies. Different predicting targets, such as $\mathbf{x}_{0}$ and $\mathbf{v}$, can also be re-parameterized to $\bm{\epsilon}$-prediction. $w(\lambda)$ denotes the loss weighting strategy. Although adjusting $w(\lambda)$ is theoretically equivalent to altering $p(\lambda)$. In practical training, directly modifying $p(\lambda)$ to concentrate computational resources on training specific noise levels is more effective than enlarging the loss weight on specific noise levels. Given these insights, our research focuses on how to design an optimal $p(\lambda)$ that can effectively allocate computational resources across different noise levels. By carefully crafting the distribution of noise intensities, we aim to improve the overall training process and the quality of the resulting diffusion models. With the unified formulation providing a flexible framework for diffusion training, we can now apply these theoretical insights to practical settings. By carefully designing the distribution of noise intensities, we can optimize the training process and improve the performance of diffusion models in real-world applications. In the following section, we will explore practical strategies for noise schedules that leverage these insights to achieve better results.
@@ 2.4 Practical Settings
By re-allocating the computation resources at different noise intensities, we can train the complete denoising process. During sampling process, we align the sampled SNRs as the cosine schedule to ensure a fair comparison. Specifically, first we sample $\{t_{0},t_{1},\ldots,t_{s}\}$ from uniform distribution $\mathcal{U}[0,1]$, then get the corresponding SNRs from Cosine schedule: $\{\frac{\alpha^{2}_{t_{0}}}{\sigma^{2}_{t_{0}}},\frac{\alpha^{2}_{t_{1}}}{\sigma^{2}_{t_{1}}},\ldots,\frac{\alpha^{2}_{t_{s}}}{\sigma^{2}_{t_{s}}}\}$. According to Equation 5, we get the corresponding $\{t^{\prime}_{0},t^{\prime}_{1},\ldots,t^{\prime}_{s}\}$ by inverting these SNR values through the respective noise schedules. Finally, we use DDIM Song et al. (2021) to sample with these new calculated $\{t^{\prime}\}$. It is important to note that, from the perspective of the noise schedule, how to allocate the computation resource during inference is also worth reconsideration. We will not explore it in this paper and leave this as future work.
@@ 3 Experiments
Dataset. We conduct experiments on ImageNet Deng et al. (2009) with $256\times 256$ and $512\times 512$ resolution. For each image, we follow the preprocessing in Rombach et al. (2022) to center crop and encode images to latents. The resulting compressed latents have dimensions of $32\times 32\times 4$ for $256^{2}$ images and $64\times 64\times 4$ for $512^{2}$ images, effectively reducing the spatial dimensions while preserving essential visual information.
Network Architecture. We adopt DiT-B from Peebles & Xie (2023) as our backbone. We replace the last AdaLN Linear layer with vanilla linear. Others are kept the same as the original implementation. The patch size is set to 2 and the projected sequence length of $32\times 32\times 4$ is $\frac{32}{2}\cdot\frac{32}{2}=256$. The class condition is injected through the adaptive layernorm. In this study, our primary objective is to demonstrate the effectiveness of our proposed noise schedule compared to existing schedules under a fixed training budget, rather than to achieve state-of-the-art results. Consequently, we do not apply our method to extra-large (XL) scale models.
Training Settings. We adopt the Adam optimizer Kingma & Ba (2014) with constant learning rate $1\times 10^{-4}$. We set the batch size to 256 following Peebles & Xie (2023) and Gao et al. (2023). Each model is trained for 500K iterations (about 100 epochs) if not specified. Our implementation is primarily based on OpenDiT Zhao et al. (2024) and experiments are mainly conducted on 8$\times$16G V100 GPUs. Different from the default discrete diffusion setting with linear noise schedule in the code base, we implement the diffusion process in a continuous way. Specifically, we sample $t$ from uniform distribution $\mathcal{U}[0,1]$.
Baselines and Metrics. We compare our proposed noise schedule with several baseline settings in Table 2. For each setting, we sample images using DDIM Song et al. (2021) with 50 steps. Despite the noise strategy for different settings may be different, we ensure they share the same $\lambda=\log\text{SNR}$ at each sampling step. This approach is adopted to exclusively investigate the impact of the noise strategy during the training phase. Moreover, we report results with different classifier-free guidance scalesHo & Salimans (2021), and the FID is calculated using 10K generated images. We sample with three CFG scales and select the optimal one to better evaluate the actual performance of different models.
@@
Table 2: Comparison of different methods and related loss weighting strategies. The $w(\lambda)$ is introduced in Equation 6. The original $p(\lambda)$ for Soft-Min-SNR Crowson et al. (2024) was developed within the EDM’s denoiser framework. In this study, we align it with the cosine schedule to ensure a fair comparison.
3.2 Comparison with baseline schedules and loss weight designs